OpenStudy (anonymous):

If four congruent circles are inscribed in a square with a side length of 20 mm as shown, find the area of the shaded region.

OpenStudy (anonymous):

I'm attaching the file...

jimthompson5910 (jim_thompson5910):

Area of one circle = 25pi Area of 4 circles = 100pi Area of square = 400 Area of shaded region = 400 - 100pi

OpenStudy (dumbcow):

Since circles are touching side of square, diameter of each circle is half side length diameter = 10 radius = 5 Area of circle = pi*5^2 = 25pi 4 circles --> total area = 100pi Area of square = 20^2 = 400 Shaded area = 400 -100pi

OpenStudy (anonymous):

400-4(25(3.14)^2)=86